OpenStudy (anonymous):

If f(x) = -1/2x -3 find f^1(x)

OpenStudy (anonymous):

f^1(x)=f(x), do you want an inverse instead?

OpenStudy (ash2326):

Lib \[f^1(x)=f(x)= \frac{-1}{2x-3}\]

OpenStudy (anonymous):

for this one i am going to guess that you have \[f(x)=-\frac{1}{2}x-3\]

OpenStudy (anonymous):

which means the inverse we find by writing \[x=-\frac{1}{2}y-3\] and solve for y \[x+3=-\frac{1}{2}y\] \[-2x-6=y\] \[f^{-1}(x)=-2x-6\]